# The gauge coupling evolutions of an ${\rm SU}(8)$ theory with the maximally symmetry breaking pattern

### v1

@@ 5.2 On the SUSY extension and the Kač-Moody extension
    in the SUSY ${\rm SU}(5)$ theory, Witten has argued that two distinctive symmetry breaking patterns of ${\rm SU}(5)\to\mathcal{G}_{\rm SM}$ and ${\rm SU}(5)\to\mathcal{G}_{41}$ are degenerate [41]. For the ${\rm SU}(8)$ theory with $\mathbf{\Phi}_{\rm Adj}\equiv\mathbf{63_{H}}$, the symmetry breaking patterns of
@@
which is quite close to the $v_{441}$ in Eq. (2). It is notable that our benchmark scales given in Ref. [10] were estimated by fitting the SM flavor parameters within the ${\rm SU}(8)$ framework, and are independent of the derivation of the string scale in Ref. [45, 46]. To achieve the string unification based on some extensions from the minimal setup of the ${\rm SU}(8)$ theory, one has to fully study all possible symmetry breaking patterns under the SUSY extensions described above, as well as to look for the constraints of the unitarity and conformal invariance from the Kač-Moody Lie algebra.
@@ Acknowledgements
We would like to thank Tianjun Li, Kaiwen Sun, Yuan Sun, Yinan Wang, Zhong-Zhi Xianyu, and Wenbin Yan for very enlightening discussions and communications. N.C. would like to thank Southeast University, and Nanjing University for hospitality when preparing this work. N.C. is partially supported by the National Natural Science Foundation of China (under Grants No. 12035008 and No. 12275140) and Nankai University. Y.N.M. is partially supported by the National Natural Science Foundation of China (under Grant No. 12205227), the Fundamental Research Funds for the Central Universities (WUT: 2022IVA052), and Wuhan University of Technology.
@@ Appendix A Conventions, decomposition rules and charge quantizations in the ${\rm SU}(8)$ theory
According to the symmetry breaking pattern in Eq. (1), we define the conventions of gauge group indices in Tab. 11. The fundamental and anti-fundamental representations will be denoted by superscripts and subscripts, respectively. We also define the decomposition rules in the ${\rm SU}(8)$ theory. After the GUT-scale symmetry breaking, we define the ${\rm U}(1)_{X_{0}}$ charges as follows
$\hat{X}_{0}(\mathbf{8})$$\equiv$ ${\rm diag}(\underbrace{-\frac{1}{4}\,,-\frac{1}{4}\,,-\frac{1}{4}\,,-\frac{1}{4}}_{\mathbf{4_{s}}}\,,\underbrace{+\frac{1}{4}\,,+\frac{1}{4}\,,+\frac{1}{4}\,,+\frac{1}{4}}_{\mathbf{4_{W}}})\,.$ (221)
@@
$\hat{X}_{1}(\mathbf{4_{s}})$$\equiv$ ${\rm diag}\,\Big{(}\underbrace{(-\frac{1}{12}+\mathcal{X}_{0})\mathbb{I}_{3\times 3}}_{\mathbf{3_{c}}}\,,\frac{1}{4}+\mathcal{X}_{0}\Big{)}\,,$ (222a)
$\hat{X}_{2}(\mathbf{4_{W}})$$\equiv$ ${\rm diag}\,\Big{(}\underbrace{(\frac{1}{12}+\mathcal{X}_{1})\mathbb{I}_{3\times 3}}_{\mathbf{3_{W}}}\,,-\frac{1}{4}+\mathcal{X}_{1}\Big{)}\,,$ (222b)
$\hat{Y}(\mathbf{4_{W}})$$\equiv$ ${\rm diag}\,\Big{(}(\frac{1}{6}+\mathcal{X}_{2})\mathbb{I}_{2\times 2}\,,-\frac{1}{3}+\mathcal{X}_{2}\,,\mathcal{X}_{2}\Big{)}$ (222c)
@@
$\hat{Q}_{e}(\mathbf{4_{W}})$$\equiv$ $T_{{\rm SU}(4)}^{3}+\hat{Y}(\mathbf{4_{W}})={\rm diag}\,\Big{(}\frac{3}{4}+\mathcal{X}_{1}\,,(-\frac{1}{4}+\mathcal{X}_{1})\mathbb{I}_{3\times 3}\Big{)}\,.$ (222d)
@@
Appendix B Note added in proof of the ${\rm SU}(8)$ maximal breaking pattern
Let us consider the maximally symmetry breaking pattern of an ${\rm SU}(N)$ group of
@@ Appendix B Note added in proof of the ${\rm SU}(8)$ maximal breaking pattern
Apparently, the condition for the odd-$N$ case has no integer solution. Thus, the unique gauge group with its unique symmetry breaking pattern where the ${\rm U}(1)$ charge does not need additional normalization factor is the ${\rm SU}(8)\to\mathcal{G}_{441}$.

### v2

@@ 5.2 On the SUSY extension and the Kač-Moody extension
    in the SUSY ${\rm SU}(5)$ theory, Witten has shown that two distinctive symmetry breaking patterns of ${\rm SU}(5)\to\mathcal{G}_{\rm SM}$ and ${\rm SU}(5)\to\mathcal{G}_{41}$ are degenerate [48]. For the ${\rm SU}(8)$ theory with $\mathbf{\Phi}_{\rm Adj}\equiv\mathbf{63_{H}}$, the symmetry breaking patterns of
@@
which is quite close to the $v_{441}$ in Eq. (2). It is notable that our benchmark scales given in Ref. [10] were estimated by fitting the SM flavor parameters within the ${\rm SU}(8)$ framework, and are independent of the derivation of the string scale in Ref. [52, 53]. To achieve the string unification based on some extensions from the minimal setup of the ${\rm SU}(8)$ theory, one has to fully study all possible symmetry breaking patterns under the SUSY extensions described above, as well as to look for the constraints of the unitarity and conformal invariance from the Kač-Moody Lie algebra [51].
@@ Acknowledgements
We would like to thank Tianjun Li, Kaiwen Sun, Yuan Sun, Yinan Wang, Zhong-Zhi Xianyu, and Wenbin Yan for very enlightening discussions and communications. N.C. would like to thank Southeast University, Nanjing University, Institute of Theoretical Physics CAS for hospitality when preparing this work. N.C. is partially supported by the National Natural Science Foundation of China (under Grants No. 12035008 and No. 12275140) and Nankai University. Y.N.M. is partially supported by the National Natural Science Foundation of China (under Grant No. 12205227), the Fundamental Research Funds for the Central Universities (WUT: 2022IVA052), and Wuhan University of Technology.
@@ Appendix A Conventions, decomposition rules and charge quantizations in the ${\rm SU}(8)$ theory
According to the symmetry breaking pattern in Eq. (1), we define the conventions of gauge group indices in Tab. 11. The fundamental and anti-fundamental representations will be denoted by superscripts and subscripts, respectively. We also define the decomposition rules in the ${\rm SU}(8)$ theory. At the zeroth GUT-scale symmetry breaking, we define the ${\rm U}(1)_{X_{0}}$ charges for the ${\rm SU}(8)$ fundamental representation as follows
$\hat{X}_{0}(\mathbf{8})$$\equiv$ ${\rm diag}(\underbrace{-\frac{1}{4}\mathbb{I}_{4\times 4}}_{\mathbf{4_{s}}}\,,\underbrace{+\frac{1}{4}\mathbb{I}_{4\times 4}}_{\mathbf{4_{W}}})\,.$ (232)
@@
${\rm first\leavevmode\nobreak\ stage}\leavevmode\nobreak$ ::: $\leavevmode\nobreak\ \hat{X}_{1}(\mathbf{4_{s}})\equiv{\rm diag}\,\Big{(}\underbrace{(-\frac{1}{12}+\mathcal{X}_{0})\mathbb{I}_{3\times 3}}_{\mathbf{3_{c}}}\,,\frac{1}{4}+\mathcal{X}_{0}\Big{)}\,,$ (233a)
${\rm second\leavevmode\nobreak\ stage}\leavevmode\nobreak$ ::: $\leavevmode\nobreak\ \hat{X}_{2}(\mathbf{4_{W}})\equiv{\rm diag}\,\Big{(}\underbrace{(\frac{1}{12}+\mathcal{X}_{1})\mathbb{I}_{3\times 3}}_{\mathbf{3_{W}}}\,,-\frac{1}{4}+\mathcal{X}_{1}\Big{)}\,,$ (233b)
${\rm third\leavevmode\nobreak\ stage}\leavevmode\nobreak$ ::: $\leavevmode\nobreak\ \hat{Y}(\mathbf{4_{W}})\equiv{\rm diag}\,\Big{(}(\frac{1}{6}+\mathcal{X}_{2})\mathbb{I}_{2\times 2}\,,-\frac{1}{3}+\mathcal{X}_{2}\,,\mathcal{X}_{2}\Big{)}$ (233c)
@@
${\rm fourth\leavevmode\nobreak\ stage}\leavevmode\nobreak$ ::: $\leavevmode\nobreak\ \hat{Q}_{e}(\mathbf{4_{W}})\equiv T_{{\rm SU}(4)}^{3}+\hat{Y}(\mathbf{4_{W}})={\rm diag}\,\Big{(}\frac{3}{4}+\mathcal{X}_{1}\,,(-\frac{1}{4}+\mathcal{X}_{1})\mathbb{I}_{3\times 3}\Big{)}\,.$ (233d)
@@
Appendix B Note added in proof of the ${\rm SU}(8)$ maximally breaking pattern
In the maximally symmetry breaking pattern of the Georgi-Glashow theory
${\rm SU}(5)\to{\rm SU}(3)_{c}\otimes{\rm SU}(2)_{W}\otimes{\rm U}(1)_{Y}\,,$ (235)
the gauge coupling of the ${\rm U}(1)_{Y}$ is normalized as
$g_{Y}{\rm diag}(-\frac{1}{3}\mathbb{I}_{3\times 3}\,,+\frac{1}{2}\mathbb{I}_{2\times 2})=g_{1}\frac{1}{\sqrt{15}}{\rm diag}(-\mathbb{I}_{3\times 3}\,,+\frac{3}{2}\mathbb{I}_{2\times 2})\Rightarrow g_{Y}^{2}=\frac{3}{5}g_{1}^{2}\,.$ (236)
Let us generalize it to the maximally symmetry breaking pattern of a generic ${\rm SU}(N)$ theory as follows
@@ Appendix B Note added in proof of the ${\rm SU}(8)$ maximally breaking pattern
Apparently, the condition for the odd-$N$ case has no integer solution. Thus, the unique gauge group with its unique maximally symmetry breaking pattern where the ${\rm U}(1)$ charge does not need additional normalization factor as in Eq. (236) is the ${\rm SU}(8)\to\mathcal{G}_{441}$. If the maximally symmetry breaking pattern was relaxed, the breaking of ${\rm SU}(8)\to\mathcal{G}_{441}$ is not the unique one where the ${\rm U}(1)$ charge does not need additional normalization. One such example is the breaking of ${\rm SU}(9)\to\mathcal{G}_{631}/\mathcal{G}_{361}$, with the ${\rm U}(1)$ charges of the ${\rm SU}(9)$ fundamental representation given by
${\rm diag}\,(-\frac{1}{6}\mathbb{I}_{6\times 6}\,,+\frac{1}{3}\mathbb{I}_{3\times 3})\,,\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm or}\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm diag}\,(-\frac{1}{3}\mathbb{I}_{3\times 3}\,,+\frac{1}{6}\mathbb{I}_{6\times 6})\,.$ (241)